In a study of obesity the following results were obtained from samples of males and females between the ages of 20 and 75:

              n         No. Overweight

Male   150           21

Female 200         48

Can we conclude from these data that in the sampled populations there is a difference in the proportions who are overweight? Let α=0.05


1
Expert's answer
2021-06-01T08:43:49-0400

Test For Significance of Different Of Proportion

Step1: Set Up Hypothesis

There Is No Significance between them - Under The Null Hypothesis "H_0: p_1 = p_2"

There Is Significance between them - Under The Alternate Hypothesis "H_1: p_1 \neq p_2"


Step2: Test Statistic

Sample 1

Probability Success "( X_1 )=21"

No. of observaed "(n_1)=150"

"P_1= X_1\/n_1=0.14"

Sample 2

Probability Success "(X_2)=48"

No. of Observaed "(n_2)=200"

"P_2= X_2\/n_2=0.24"


Finding a "\hat P" value For Proportion "\hat P=\dfrac{X_1+X_2}{n_1+n_2}=0.1971"


"\hat Q" Value For Proportion= "1-\hat P=0.8029"


if n>30 So, we use Test Statistic "(Z) = \dfrac{(P_1-P_2)}{(\sqrt{\hat P\cdot \hat Q(\frac{1}{n_1}+\frac{1}{n_2}}))}=\dfrac{0.14-0.24}{\sqrt{0.197\times 0.8029(\frac{1}{150}+\frac{1}{200})}}"



"Z_{ cal}=-2.32711\\

| Z _{cal} | =2.3271"


Step3: Tabulated Value

The Value of "|Z_{ tab}|" at LOS 0.05% is 1.96

We got "|Z _{cal}| =2.3271" & "| Z_{ tab} | =1.96"


Step4: Make Decision

Hence Value of "| Z_{ cal} | > | Z _{tab}|" and Here we Accept Ho

There Is Significance between them, i.e there is a difference in the proportions


 P-Value

Two Tailed ( double the one tail ):

"H_1 : ( P \neq -2.3271 ) = 0.02"